3x+3y=18
3x-5y=2
What is x and what is y???

Answer by stanbon(75887)

3x+3y=18
3x-5y=2
----
Subtract and solve for "y":
8y = 16
y = 2
------------
Solve for "x":
3x-5y = 2
3x -5*2 = 2
3x = 12
x = 4
-------
Solution: (4,2)
================
Cheers,
Stan H.
================